OpenStudy (anonymous):

if it takes tyler 2 hours to do a project and it takes his friend zach 4 hours to do the same project, how long would it take them to do the same project together?

OpenStudy (anonymous):

igot 3hours

OpenStudy (anonymous):

I Will Take Them 6 hOURS TO DO THE PROJECT

OpenStudy (anonymous):

total time 1/t=1/2+1/4 1/t= 6/8 t=8/6 hours

OpenStudy (anonymous):

can you explain how